# HAODiff: Human-Aware One-Step Diffusion via Dual-Prompt Guidance

### v1

@@ 3.1 Degradation Pipeline with Human Motion Blur (HMB)
Figure 4: Structure of the prompt embedder. The Attention Pooling uses a learnable embedding as $Q$, while $K$ and $V$ from the output of Performer Encoder, whose depth $N$ is set to 6.
As for the prompt extractor part, after training the dual-prompt guidance (DPG) in Sec. 3.2, we further adapt it to be compatible with providing prompt embeddings. We extract the output feature from the last layer of the $H_{Ri}$ to obtain embedding features with the most informative representation. However, the size of this output significantly differs from the embedding size required by SD, and its feature space distribution also deviates from standard text embeddings. Therefore, before integrating into Stable Diffusion (SD), we apply a nonlinear mapping and feature compression using a linear complexity Performer choromanski2021performer Encoder and an attention pooling lee2019attentionpool module, denoted prompt embedder, as shown in Fig. 4. The features of the two negative branches are concatenated and, with the positive branch, fed into different prompt embedders. This process yields dual-prompt embeddings $p_{\text{pos}}$ and $p_{\text{neg}}$. Finally, by concatenating these embeddings along the batch size dimension, the receiving UNet ronneberger2015unet can efficiently obtain both $z_{\text{pos}}$ and $z_{\text{neg}}$ in parallel.
Training Objective. The human body restoration model aims to recover high-quality human images with rich details from degraded inputs. During training, we use pixel-level mean squared error loss $\mathcal{L}_{\text{MSE}}$ to minimize reconstruction errors. To enhance edge responses, we also employ an edge-aware DISTS perceptual loss, denoted as $\mathcal{L}_{\text{EA}}$. It is calculated by feeding both the original image and its edge-enhanced version, obtained using the Sobel operator $\mathcal{S}(\cdot)$, into the DISTS ding2020dists function:
@@
Previous study 2024s3diff shows that even when the restored image is similar to the original, distortions may still occur in the latent vector distribution. To address this, we utilize a pre-trained SD UNet downsampling module $\mathcal{D}_{\psi}$ as a discriminator and calculate the generator loss $\mathcal{L}_{\mathcal{G}}$. It helps the model learn a more realistic data distribution. The total loss function $\mathcal{L}_{\text{total}}$ is defined as follows:
@@
The generator discriminator network (GAN) goodfellow2014gan loss consists of the generator loss $\mathcal{L}_{\mathcal{G}}$ and the discriminator loss $\mathcal{L}_{\mathcal{D}}$. Following previous works 2024s3diff ; wang2024osdface , we define them as follows:
@@
SUPIR yu2024supir 50 26.67 0.1415 0.2929 0.4080 13.8357 0.7908 0.6811 3.7769 4.2303 21.3221
DiffBIR lin2024diffbir 50 9.03 0.1402 0.2797 0.4319 12.9306 0.7792 0.6965 3.9404 4.4211 21.4859
SeeSR wu2024seesr 50 5.05 0.1295 0.2555 0.4476 12.8225 0.7620 0.6979 3.5744 4.5722 21.1563
PASD yang2023pasd 20 3.15 0.1469 0.2910 0.4383 15.6758 0.6121 0.6477 3.8042 4.0878 25.4919
ResShift yue2023resshift 15 2.88 0.1638 0.2848 0.4237 15.6502 0.5365 0.5650 5.2973 3.3929 25.3706
SinSR wang2024sinsr 1 0.19 0.1579 0.2840 0.4037 16.9203 0.6037 0.5802 4.4221 3.6340 21.1563
OSEDiff wu2024osediff 1 0.13 0.1340 0.2507 0.4539 14.7489 0.6961 0.6769 3.4689 4.7567 21.9784
InvSR yue2024invsr 1 0.17 0.1424 0.2709 0.4297 13.2950 0.6886 0.6845 3.7446 4.3290 21.6265
OSDHuman gong2025osdhuman 1 0.11 0.1356 0.2384 0.4680 14.4121 0.7312 0.6908 3.8278 4.7512 22.3524
@@
Table 1: Quantitative results on PERSONA-Val and inference time comparison. The top two scores are highlighted in red and cyan for all methods. C-IQA stands for CLIPIQA, and M-IQA stands for MANIQA. Inference is performed on images of 512$\times$512 resolution on NVIDIA RTX A6000.
@@ 4.1 Experimental Settings
Training and Testing Datasets. Our model is trained on the PERSONA gong2025osdhuman , with 20k images sampled from both LSDIR Li2023LSDIR and FFHQ karras2019ffhq . We randomly crop LSDIR to 512$\times$512, and pre-downsample FFHQ to the same size. We generate synthetic HQ-LQ image pairs using our degradation pipeline. For testing, we use PERSONA-Val and PERSONA-Test from OSDHuman gong2025osdhuman . Additionally, we select images from the MPII Human Pose dataset andriluka2014mpii using the data selection pipeline from OSDHuman, excluding the quality filtering stage. To accommodate the bounding boxes used in this process, we extend the outermost annotated key points outward by a certain margin to form enclosing rectangles, which are used as bounding boxes. This process yields MPII-Test, which consists of 5,427 real-world images with diverse human motion blur (HMB). Using our degradation pipeline, we fine-tune YOLO11 yolo11_ultralytics to detect HMB, finding 1,765 instances in 1,326 images.
Evaluation Metrics. For PERSONA-Val, we use both full-reference and no-reference metrics. For full-reference perceptual quality assessment, we use DISTS ding2020dists , LPIPS zhang2018lpips , and TOPIQ Chen2024topiq . Additionally, we calculate FID heusel2017fid to measure the distribution distance between the restored images and ground truth. For no-reference metrics, we use CLIPIQA wang2022clipiqa , NIQE zhang2015niqe , MANIQA-pipal yang2022maniqa , LIQE zhang2023liqe , and IL-NIQE Zhang2015ilniqe . These no-reference metrics are also applied to PERSONA-Test and MPII-Test. Additionally, for MPII-Test, we use the fine-tuned YOLO model to detect HMB instances and calculate the ratio of detected instances in restored images to those in original images, denoted as HMB-R. The details of YOLO in HMB detection are discussed in the supplementary material.
Implementation Details. In stage 1, the DPG training process balances the magnitude of the L1 loss and Dice loss by setting the $\alpha$ in Eq. (4) to 2$\times$10-2. We use the Adam optimizer kingma2017adam with learning rate 2$\times$10-3 and batch size 16. The STLs in $H_{E}$ have 6 heads, while those in $H_{Ri}$ use 3 heads. For HBR segmentation, the third branch outputs a single channel and utilizes the sigmoid activation function, while the other two output three channels without the activation function. The training is conducted for 20k iterations on 4 NVIDIA RTX A6000 GPUs. In stage 2, we set $\beta$ in Eq. (10) to 1×1$\times$1 ×10-2 and use the pretrained SDXL podell2023sdxl UNet as the discriminator following D3SR li2025d3sr . The $\lambda_{\text{cfg}}$ in Eq. (8) is set to 3.5. The AdamW optimizer loshchilov2018AdamW used in stage 2 has learning rate 1$\times$10-5 and batch size 2. The base model is SD2.1-base sd21 and LoRA hu2022lora is used to train the UNet with a LoRA rank 16. The training is conducted for 120k iterations on 2 NVIDIA RTX A6000 GPUs.
Compared State-of-the-Art (SOTA) Methods. We compare HAODiff with multi-step diffusion models, including DiffBIR lin2024diffbir , SeeSR wu2024seesr , SUPIR yu2024supir , PASD yang2023pasd and ResShift yue2023resshift ; and one-step diffusion models, including SinSR wang2024sinsr , OSEDiff wu2024osediff , InvSR yue2024invsr , and OSDHuman gong2025osdhuman . Comparisons with other restoration and deblurring methods are shown in the supplementary material.
@@
SUPIR yu2024supir 0.7106 0.6798 4.0734 4.0501 22.6182 0.6702 0.6256 4.4231 3.4295 26.0135 0.2776
DiffBIR lin2024diffbir 0.7287 0.6812 4.9820 3.8870 25.4183 0.6531 0.6405 5.1638 3.1343 27.8964 0.5705
SeeSR wu2024seesr 0.6968 0.6759 4.1126 4.0800 23.4376 0.6478 0.6636 4.6152 3.5709 26.5874 0.2612
PASD yang2023pasd 0.5765 0.6703 3.8728 3.8901 24.9226 0.4023 0.5220 6.0325 2.3903 38.3461 0.9518
ResShift yue2023resshift 0.5544 0.6101 4.8438 3.4981 25.0764 0.4356 0.5128 6.4687 2.4915 32.7579 1.0799
SinSR wang2024sinsr 0.5882 0.6010 4.7510 3.5339 23.3862 0.4816 0.5162 5.9015 2.4504 29.0120 0.9394
OSEDiff wu2024osediff 0.6734 0.6919 4.4600 4.4296 24.4183 0.6385 0.6580 4.8383 4.1664 26.8298 0.1853
InvSR yue2024invsr 0.6837 0.7122 4.1694 4.2582 22.8932 0.6166 0.6573 5.4187 3.5793 26.8856 0.2771
OSDHuman gong2025osdhuman 0.7155 0.6977 4.1287 4.3202 24.8712 0.6537 0.6471 4.5960 3.7991 26.8940 0.3433
@@
Table 2: Quantitative comparison on real and synthetic benchmarks, with top two results respectively highlighted in red and cyan. HMB-R indicates the ratio of human motion blur instances compared with restored and original images. C-IQA stands for CLIPIQA, and M-IQA stands for MANIQA.
@@ 4.2 Main Results
Quantitative Comparisons. Table 1 presents quantitative results on the synthesized PERSONA-Val. Despite HAODiff’s overwhelming inference speed advantage over multi-step diffusion models and comparable performance to some one-step diffusion (OSD) models, it achieves top scores across all full-reference metrics. In no-reference evaluation, it ranks first in NIQE and IL-NIQE, indicating natural image synthesis, and leads in MANIQA and LIQE, reflecting alignment with human aesthetics. Among OSD models, HAODiff exhibits the best performance in CLIPIQA, showcasing high-quality restoration. Table 2 compares HAODiff on PERSONA-Test and MPII-Test. On PERSONA-Test, HAODiff maintains superior performance, leading on most metrics. For the human motion blur (HMB)-rich MPII-Test, HAODiff tops every metric. Meanwhile, our model achieved the lowest HMB-R, demonstrating strong HBR performance and effective restoration of HMB content.
Qualitative Comparisons. Visual comparisons with SOTA methods are shown in Figs. 5 and 6. HAODiff effectively restores generic degradations in low-quality (LQ) images, producing visually clear results. On the LQ human body dataset PERSONA-Test, our model achieves notably better reconstruction of both body and facial regions. Specifically, it is capable of restoring the applied HMB. On PERSONA-Val and MPII-Test, HAODiff shows clear advantages in handling human images with prominent HMB artifacts, delivering effective restoration. In regions affected by severe motion blur, the model produces results with sharp contours and realistic visual quality. Compared to both one-step and multi-step diffusion models, HAODiff consistently generates more detailed and natural results. It avoids the excessive smoothness observed in PASD yang2023pasd , ResShift yue2023resshift , and OSEDiff wu2024osediff , as well as the over-sharpened and distorted textures found in SUPIR yu2024supir and InvSR yue2024invsr . In addition, HAODiff produces more realistic clothing details, accurately reflecting fabric textures and folds under natural lighting conditions. Moreover, HAODiff specifically restores blurred human regions without altering the depth of field, conforming to natural optical factors, thereby preventing unrealistic texture synthesis in the background. Further visual comparisons are provided in the supplementary material.
@@ 4.3 Ablation Study
HMB-aware Branch of DPG. To assess the impact of the human motion blur (HMB)-aware branch (the third branch) in addressing HMB, we conducted a targeted ablation study. Specifically, we train HAODiff using a version of DPG that excludes the HMB-aware branch while retaining the other two branches. The training is performed for the same number of epochs, with identical settings and CFG strategy, and the model is further evaluated on the HMB-rich MPII-Test. The results are shown in Fig. 7. Without the guidance of the HMB-aware branch, the model shows a clear decline in its ability to recover the HMB regions. Specifically, the arms of the subject in the LQ image suffer from severe motion blur; without the branch, the model removes only non-motion-blur noise, whereas with the branch the arms are sharply restored. Quantitative comparisons between the two configurations are also provided, with LIQE and HMB-R scores annotated below each strategy. These results indicate that incorporating the HMB-aware branch not only preserves the overall restoration quality but also significantly improves the model’s ability to handle HMB-specific degradation.
@@
Text Pair 2024s3diff 0.1056 0.3815 8.5215 0.7648 2.9692
DAPE wu2024osediff 0.1047 0.4296 8.6143 0.7693 3.0779
HFIE gong2025osdhuman 0.1060 0.4066 9.2570 0.7681 3.2063
@@
The Effectiveness of DPG. In order to validate the effectiveness of DPG, we compare it against several prompt-based guidance strategies, including: the fixed positive–negative prompt pair with CFG following S3Diff 2024s3diff , the degradation aware prompt extractor (DAPE) from OSEDiff wu2024osediff , and the high-fidelity image embedder (HFIE) from OSDHuman gong2025osdhuman . Table 3 reports quantitative results on PERSONA-Val, where DPG consistently delivers superior restoration quality across multiple metrics. Moreover, we observe that other methods tend to produce more pronounced color shifts compared to DPG-guided outputs. These color shifts significantly affect human perception of image fidelity. And once they occur, it is challenging to eliminate them through post-processing methods (e.g., wavelet-based correction mallat1989wavelet ). To quantify this effect and assess the impact on model stability, we employ MS-SWD he2024msswd , a metric for measuring color distribution differences between the processed results and reference images. With HQ images as reference, the findings clearly show that DPG yields the smallest color discrepancies. DPG’s superior performance in this respect stems from its dual-prompt design: the negative path captures residual noise, implicitly including color-shift degradation. During training, DPG more accurately anchors the original color distribution and detects color drift in the noise. This capability is then explicitly reinforced through classifier-free guidance (CFG), thereby resulting in higher fidelity and robustness.

### v2

@@ 3.1 Degradation Pipeline with Human Motion Blur (HMB)
Figure 4: Structure of the prompt embedder. The Attention Pooling uses a learnable embedding as $Q$, while $K$ and $V$ from the output of Performer Encoder, whose depth NN is set to 6.
As for the prompt extractor part, after training the dual-prompt guidance (DPG) in Sec. 3.2, we further adapt it to be compatible with providing prompt embeddings. We extract the output feature from the last layer of the $H_{Ri}$ to obtain embedding features with the most informative representation. However, the size of this output significantly differs from the embedding size required by SD, and its feature space distribution also deviates from standard text embeddings. Therefore, before integrating into Stable Diffusion (SD), we apply a nonlinear mapping and feature compression using a linear complexity Performer Choromanski et al. (2021) Encoder and an attention pooling Lee et al. (2019) module, denoted prompt embedder, as shown in Fig. 4. The features of the two negative branches are concatenated and, with the positive branch, fed into different prompt embedders. This process yields dual-prompt embeddings $p_{\text{pos}}$ and $p_{\text{neg}}$. Finally, by concatenating these embeddings along the batch size dimension, the receiving UNet Ronneberger et al. (2015) can efficiently obtain both $z_{\text{pos}}$ and $z_{\text{neg}}$ in parallel.
Training Objective. The human body restoration model aims to recover high-quality human images with rich details from degraded inputs. During training, we use pixel-level mean squared error loss $\mathcal{L}_{\text{MSE}}$ to minimize reconstruction errors. To enhance edge responses, we also employ an edge-aware DISTS perceptual loss, denoted as $\mathcal{L}_{\text{EA}}$. It is calculated by feeding both the original image and its edge-enhanced version, obtained using the Sobel operator $\mathcal{S}(\cdot)$, into the DISTS Ding et al. (2020) function:
@@
Previous study Zhang et al. (2024a) shows that even when the restored image is similar to the original, distortions may still occur in the latent vector distribution. To address this, we utilize a pre-trained SD UNet downsampling module $\mathcal{D}_{\psi}$ as a discriminator and calculate the generator loss $\mathcal{L}_{\mathcal{G}}$. It helps the model learn a more realistic data distribution. The total loss function $\mathcal{L}_{\text{total}}$ is defined as follows:
@@
The generative adversarial network (GAN) Goodfellow et al. (2014) loss consists of the generator loss $\mathcal{L}_{\mathcal{G}}$ and the discriminator loss $\mathcal{L}_{\mathcal{D}}$. Following previous works Zhang et al. (2024a); Wang et al. (2025), we define them as follows:
@@
SUPIR Yu et al. (2024) 50 26.67 0.1415 0.2929 0.4080 13.8357 0.7908 0.6811 3.7769 4.2303 21.3221
DiffBIR Lin et al. (2024) 50 9.03 0.1402 0.2797 0.4319 12.9306 0.7792 0.6965 3.9404 4.4211 21.4859
SeeSR Wu et al. (2024b) 50 5.05 0.1295 0.2555 0.4476 12.8225 0.7620 0.6979 3.5744 4.5722 21.1563
PASD Yang et al. (2023) 20 3.15 0.1469 0.2910 0.4383 15.6758 0.6121 0.6477 3.8042 4.0878 25.4919
ResShift Yue et al. (2023) 15 2.88 0.1638 0.2848 0.4237 15.6502 0.5365 0.5650 5.2973 3.3929 25.3706
SinSR Wang et al. (2024c) 1 0.19 0.1579 0.2840 0.4037 16.9203 0.6037 0.5802 4.4221 3.6340 21.1563
OSEDiff Wu et al. (2024a) 1 0.13 0.1340 0.2507 0.4539 14.7489 0.6961 0.6769 3.4689 4.7567 21.9784
InvSR Yue et al. (2025) 1 0.17 0.1424 0.2709 0.4297 13.2950 0.6886 0.6845 3.7446 4.3290 21.6265
OSDHuman Gong et al. (2025) 1 0.11 0.1356 0.2384 0.4680 14.4121 0.7312 0.6908 3.8278 4.7512 22.3524
@@
Table 1: Quantitative results on PERSONA-Val and inference time comparison. The top two scores are highlighted in red and cyan for all methods. C-IQA stands for CLIPIQA, and M-IQA stands for MANIQA. Inference is performed on images of $\times512$ resolution on NVIDIA RTX A6000.
@@ 4.1 Experimental Settings
Training and Testing Datasets. Our model is trained on the PERSONA Gong et al. (2025), with 20k images sampled from both LSDIR Li et al. (2023b) and FFHQ Karras et al. (2019). We randomly crop LSDIR to $\times512$, and pre-downsample FFHQ to the same size. We generate synthetic HQ-LQ image pairs using our degradation pipeline. For testing, we use PERSONA-Val and PERSONA-Test from OSDHuman Gong et al. (2025). Additionally, we select images from the MPII Human Pose dataset Andriluka et al. (2014) using the data selection pipeline from OSDHuman, excluding the quality filtering stage. To accommodate the bounding boxes used in this process, we extend the outermost annotated key points outward by a certain margin to form enclosing rectangles, which are used as bounding boxes. This process yields MPII-Test, which consists of 5,427 real-world images with diverse human motion blur (HMB). Using our degradation pipeline, we fine-tune YOLO11 Jocher and Qiu (2024) to detect HMB, finding 1,765 instances in 1,326 images.
Evaluation Metrics. For PERSONA-Val, we use both full-reference and no-reference metrics. For full-reference perceptual quality assessment, we use DISTS Ding et al. (2020), LPIPS Zhang et al. (2018), and TOPIQ Chen et al. (2024). Additionally, we calculate FID Heusel et al. (2017) to measure the distribution distance between the restored images and ground truth. For no-reference metrics, we use CLIPIQA Wang et al. (2023), NIQE Zhang et al. (2015b), MANIQA-pipal Yang et al. (2022), LIQE Zhang et al. (2023), and IL-NIQE Zhang et al. (2015a). These no-reference metrics are also applied to PERSONA-Test and MPII-Test. Additionally, for MPII-Test, we use the fine-tuned YOLO model to detect HMB instances and calculate the ratio of detected instances in restored images to those in original images, denoted as HMB-R. The fine-tuned YOLO detector is evaluated on a dedicated hold-out set containing 4,216 images, achieving an mAP@0.5 of 0.6223 and a precision of 0.6738.
[…]
Compared State-of-the-Art (SOTA) Methods. We compare HAODiff with multi-step diffusion models, including DiffBIR Lin et al. (2024), SeeSR Wu et al. (2024b), SUPIR Yu et al. (2024), PASD Yang et al. (2023) and ResShift Yue et al. (2023); and one-step diffusion models, including SinSR Wang et al. (2024c), OSEDiff Wu et al. (2024a), InvSR Yue et al. (2025), and OSDHuman Gong et al. (2025).
@@
SUPIR Yu et al. (2024) 0.7106 0.6798 4.0734 4.0501 22.6182 0.6702 0.6256 4.4231 3.4295 26.0135 0.2776
DiffBIR Lin et al. (2024) 0.7287 0.6812 4.9820 3.8870 25.4183 0.6531 0.6405 5.1638 3.1343 27.8964 0.5705
SeeSR Wu et al. (2024b) 0.6968 0.6759 4.1126 4.0800 23.4376 0.6478 0.6636 4.6152 3.5709 26.5874 0.2612
PASD Yang et al. (2023) 0.5765 0.6703 3.8728 3.8901 24.9226 0.4023 0.5220 6.0325 2.3903 38.3461 0.9518
ResShift Yue et al. (2023) 0.5544 0.6101 4.8438 3.4981 25.0764 0.4356 0.5128 6.4687 2.4915 32.7579 1.0799
SinSR Wang et al. (2024c) 0.5882 0.6010 4.7510 3.5339 23.3862 0.4816 0.5162 5.9015 2.4504 29.0120 0.9394
OSEDiff Wu et al. (2024a) 0.6734 0.6919 4.4600 4.4296 24.4183 0.6385 0.6580 4.8383 4.1664 26.8298 0.1853
InvSR Yue et al. (2025) 0.6837 0.7122 4.1694 4.2582 22.8932 0.6166 0.6573 5.4187 3.5793 26.8856 0.2771
OSDHuman Gong et al. (2025) 0.7155 0.6977 4.1287 4.3202 24.8712 0.6537 0.6471 4.5960 3.7991 26.8940 0.3433
@@
Table 2: Quantitative comparison on real-world benchmarks, with top two results respectively highlighted in red and cyan. HMB-R indicates the ratio of human motion blur instances compared with restored and original images. C-IQA stands for CLIPIQA, and M-IQA stands for MANIQA.
@@ 4.2 Main Results
Quantitative Comparisons. Table 1 presents quantitative results on the synthesized PERSONA-Val. Despite HAODiff’s overwhelming inference speed advantage over multi-step diffusion models and comparable performance to some one-step diffusion (OSD) models, it achieves top scores across all full-reference metrics. In no-reference evaluation, it ranks first in NIQE and IL-NIQE, indicating natural image synthesis, and leads in MANIQA and LIQE, reflecting alignment with human aesthetics. Among OSD models, HAODiff exhibits the best performance in CLIPIQA, showcasing high-quality restoration. Table 2 compares HAODiff on PERSONA‑Test and MPII‑Test. On PERSONA-Test, HAODiff maintains superior performance, leading on most metrics. For the human motion blur (HMB)‑rich MPII‑Test, HAODiff tops every metric. Meanwhile, our model achieved the lowest HMB-R, demonstrating strong HBR performance and effective restoration of HMB content.
Qualitative Comparisons. Visual comparisons with SOTA methods are shown in Figs. 5 and 6. HAODiff effectively restores generic degradations in low-quality (LQ) images, producing visually clear results. On the LQ human body dataset PERSONA-Test, our model achieves notably better reconstruction of both body and facial regions. Specifically, it is capable of restoring the applied HMB. On PERSONA-Val and MPII-Test, HAODiff shows clear advantages in handling human images with prominent HMB artifacts, delivering effective restoration. In regions affected by severe motion blur, the model produces results with sharp contours and realistic visual quality. Compared to both one-step and multi-step diffusion models, HAODiff consistently generates more detailed and natural results. It avoids the excessive smoothness observed in PASD Yang et al. (2023), ResShift Yue et al. (2023), and OSEDiff Wu et al. (2024a), as well as the over-sharpened and distorted textures found in SUPIR Yu et al. (2024) and InvSR Yue et al. (2025). In addition, HAODiff produces more realistic clothing details, accurately reflecting fabric textures and folds under natural lighting conditions. Moreover, HAODiff specifically restores blurred human regions without altering the depth of field, conforming to natural optical factors, thereby preventing unrealistic texture synthesis in the background. Further visual comparisons are provided in the supplementary material.
@@ 4.3 Ablation Study
HMB-aware Branch of DPG. To assess the impact of the human motion blur (HMB)-aware branch (the third branch) in addressing HMB, we conducted a targeted ablation study. Specifically, we train HAODiff using a version of DPG that excludes the HMB-aware branch while retaining the other two branches. The training is performed for the same number of epochs, with identical settings and CFG strategy, and the model is further evaluated on the HMB‑rich MPII‑Test. The results are shown in Fig. 7. Without the guidance of the HMB-aware branch, the model shows a clear decline in its ability to recover the HMB regions. Specifically, the arms of the subject in the LQ image suffer from severe motion blur; without the branch, the model removes only non‑motion-blur noise, whereas with the branch the arms are sharply restored. Quantitative comparisons between the two configurations are also provided, with LIQE and HMB-R scores annotated below each strategy. These results indicate that incorporating the HMB-aware branch not only preserves the overall restoration quality but also significantly improves the model’s ability to handle HMB-specific degradation.
@@
Text Pair Zhang et al. (2024a) 0.1056 0.3815 8.5215 0.7648 2.9692
DAPE Wu et al. (2024a) 0.1047 0.4296 8.6143 0.7693 3.0779
HFIE Gong et al. (2025) 0.1060 0.4066 9.2570 0.7681 3.2063
@@
The Effectiveness of DPG. In order to validate the effectiveness of DPG, we compare it against several prompt-based guidance strategies, including: the fixed positive–negative prompt pair with CFG following S3Diff Zhang et al. (2024a), the degradation aware prompt extractor (DAPE) from OSEDiff Wu et al. (2024a), and the high-fidelity image embedder (HFIE) from OSDHuman Gong et al. (2025). Table 3 reports quantitative results on PERSONA-Val, where DPG consistently delivers superior restoration quality across multiple metrics. Moreover, we observe that other methods tend to produce more pronounced color shifts compared to DPG-guided outputs. These color shifts significantly affect human perception of image fidelity. And once they occur, it is challenging to eliminate them through post-processing methods (e.g., wavelet-based correction Mallat (1989)). To quantify this effect and assess the impact on model stability, we employ MS‑SWD He et al. (2024), a metric for measuring color distribution differences between the processed results and reference images. With HQ images as reference, the findings clearly show that DPG yields the smallest color discrepancies. DPG’s superior performance in this respect stems from its dual-prompt design: the negative path captures residual noise, implicitly including color-shift degradation. During training, DPG more accurately anchors the original color distribution and detects color drift in the noise. This capability is then explicitly reinforced through classifier-free guidance (CFG), thereby resulting in higher fidelity and robustness.